OpenStudy (anonymous):

Solve log 2x2 + 2 = 6. Round to the nearest thousandth if necessary.

Nnesha (nnesha):

is it \[\huge\rm log 2x^2 +2 =6\] ?

OpenStudy (anonymous):

I'm getting \[50\sqrt{2} \] or just about 70.711

Nnesha (nnesha):

is it \[\huge\rm log (2x^2 +2 )=6\] ? or 2 is base ?

OpenStudy (trwatkins1):

can help :( ???

OpenStudy (trwatkins1):

*cant

OpenStudy (anonymous):

its um log 2x^2 + 2 = 6 (log and 2x^2 are seperate)

Nnesha (nnesha):

oh so this is right \[\huge\rm log (2x^2 +2 )=6\] 2 isn't base

Nnesha (nnesha):

alright can you please explain how you got the answer :=) what was ur first step ? :=)

Nnesha (nnesha):

hmm \[\huge\rm log_2 (x^2+2)=6\] \[\huge\rm log (2x^2+2)=6\] which is ur question ?

OpenStudy (anonymous):

it's the second one

OpenStudy (anonymous):

I subtract 2 from both sides log10 (2x^2) + 2 -2 = 6-2 log10 (2x^2)=4 I then used a = logb (b^a) so... 4= log10 (10^4) log10 (2x^2) = log10 (10^4) then all this other stuff/ not sure I want to type it all out lol

OpenStudy (anonymous):

@Nnesha

OpenStudy (triciaal):